# Negative Sampling in Recommendation: A Survey and Future Directions

### v1

@@ Abstract.
Recommender systems aim to capture users’ personalized preferences from the cast amount of user behaviors, making them pivotal in the era of information explosion. However, the presence of the dynamic preference, the “information cocoons”, and the inherent feedback loops in recommendation make users interact with a limited number of items. Conventional recommendation algorithms typically focus on the positive historical behaviors, while neglecting the essential role of negative feedback in user interest understanding. As a promising but easy-to-ignored area, negative sampling is proficients in revealing the genuine negative aspect inherent in user behaviors, emerging as an inescapable procedure in recommendation. In this survey, we first discuss the role of negative sampling in recommendation and thoroughly analyze challenges that consistently impede its progress. Then, we conduct an extensive literature review on the existing negative sampling strategies in recommendation and classify them into five categories with their discrepant techniques. Finally, we detail the insights of the tailored negative sampling strategies in diverse recommendation scenarios and outline an overview of the prospective research directions toward which the community may engage and benefit.
@@ 1. Introduction
In alignment with other supervised tasks, providing appropriate supervised signals (both positive and negative feedback) is indispensable in the training phase of recommendation algorithms. The absence of either of these feedback types can inevitably give rise to bias in model training (Ma et al., 2023; Zhu et al., 2022; Chen et al., 2023b). However, as illustrated in Fig. 1, explicit feedback remains conspicuously absent within most recommendation datasets. Even when encompassed within datasets of such information, certain studies (Zhang et al., 2013; Ding et al., 2020; Chen et al., 2023b) have unearthed that these negative feedback indeed mirror user preferences to some extent. For instance, the behavior of assigning a low rating to a particular movie may paradoxically signify that the user has already viewed the movie, as this selection from an extensive movie pool subtly reveals the user’s underlying predilections (Ding et al., 2023; Huang et al., 2023). Consequently, the integral exploration of negative sampling strategies that are proficient in unveiling genuine negative aspects inherent in user preferences emerges as an inescapable procedure in RS.
@@
The profound impact of distinct negative sampling strategies on the performance of recommendation algorithms in disparate recommendation scenarios underscores the intricacies inherent in this critical domain. Moreover, existing academic endeavors pertaining to negative sampling algorithms are diverse in RS. Varied resolutions are observed in addressing negative samples across distinct recommendation scenarios, indicating the absence of a universal algorithm capable of concurrently tackling all recommendation tasks (Cheng et al., 2021; Li et al., 2021; Liu et al., 2021b; Zhao et al., 2019; Shi et al., 2023; Gong et al., 2022; Zhang et al., 2013; Kim et al., 2021a; Yang et al., 2022a). Finally, the predominant strategies for tackling matching and ranking tasks revolve around the random sampling strategies (Zhao et al., 2016) and the meticulous integration of genuine negative feedback (Lian et al., 2020a; Shen et al., 2021; Gu et al., 2021) for the industrial recommendation systems. Currently, there lacks a comprehensive survey that systematically categorizes and presents the existing studies. Such a survey can serve as a handbook to offer a unified exploration of the varied landscape surrounding negative sampling algorithms in RS, thereby addressing the aforementioned issues.
@@
Given the significance and popularity of RS, there are several recently published surveys also reviewed this area from different perspectives, including debiased recommendation (Chen et al., 2023a), cross-domain recommendation (Zang et al., 2022), graph-based recommendation (Gao et al., 2023c), knowledge graph-based recommendation (Guo et al., 2020b), causal inference-based recommendation (Gao et al., 2022b), large language model-based recommendation (Fan et al., 2023; Wu et al., 2023) and evaluation in recommendation (Bauer et al., 2023). As to negative sampling, there are surveys in other domains, including knowledge graph completion (Cai et al., 2022), deep metric learning (Kaya and Bilge, 2019), neural semantic retrieval (Guo et al., 2022; Chen et al., 2022c), image retrieval (Chen et al., 2022c), and representation learning on various types of data (Liu and Tang, 2021; Qian et al., 2020; Xu et al., 2022; Yang et al., 2020b; Chen et al., 2020a; Kamigaito and Hayashi, 2022; Wang et al., 2021a; Khoshraftar and An, 2022; Wang et al., 2017a). Compared to these research domains, negative sampling techniques employed in RS possess distinctive characteristics, necessitating a comprehensive review and summary. To bridge this gap, we review the negative sampling strategies in RS, classify them into six categories with the discrepant techniques and detail the insights of the tailored negative sampling strategies in diverse recommendation scenarios.
@@
This survey encompasses over 180 papers from the top conferences and journals within the field of recommendation systems, which focus on the analysis of negative samples and the design of novel negative sampling techniques. We provide a comprehensive evaluation of negative sampling strategies in recommendation from various perspectives. In particular, we search related works from the top-tier conferences in recommendation, including SIGIR, WWW, KDD, ACL, AAAI, IJCAI, ICDM, ICDE, RecSys, WSDM, CIKM, and PAKDD, as well as the leading journals such as TOIS, TKDE, TKDD, KBS, TMM, etc, with the keywords ”recommendation”, ”collaborative filtering”, ”ranking” and ”matching” in conjunction with ”negative samples”, ”re-weighting” and ”sampling”. At the same time, to prevent omissions of relevant literature, we utilize Google Scholar, DBLP, and Arxiv to search for related studies. After getting the list of papers, we also check their references to retain the unobserved papers about negative sampling strategies. To facilitate this survey’s role as a handbook for negative sampling techniques in RS, we will release a GitHub repository to encompass all reviewed papers along with their corresponding codes at the following link: https://github.com/hulkima/NS4RS.
@@ 2.1. Role of Negative Sampling in Recommendation
Real-world recommendation systems frequently involve more than millions of users and items, rendering the integration of all corpus into the training process prohibitively expensive (Zhang et al., 2013; Shi et al., 2023; Zhou et al., 2020). Despite the presence of a substantial number of users and items, the existence of “Information Cocoons” and the unique feedback loops within recommendation result in the majority of users interacting with only a limited number of items in the corpus (Li et al., 2022; Lou et al., 2019; Chen et al., 2021c). As a consequence, this gives rise to the well-documented issue of data sparsity in recommendations (Ma et al., 2024a, 2023; Cao et al., 2022; Xie et al., 2022b). The dynamic nature of user interests also necessitates a continuous process of adaptation and updating for recommenders (Zhao et al., 2016; Xu et al., 2021; He et al., 2016). Furthermore, the perpetual introduction of new users and items (both devoid of an adequate reservoir of historical interaction data) engenders the notorious cold start problem (Wei et al., 2021; Togashi et al., 2021).
@@
Large-scale recommendation applications typically encompass multiple phases to narrow down the candidate items for the recommender, thereby alleviating the aforementioned issues and significantly reducing their computational complexity (Huang et al., 2023; Ding et al., 2023). These algorithms generally require both positive and negative examples to model the users’ personalized preferences. Negative sampling is the critical and irreplaceable element in recommendation that could potentially improve the modeling of dynamic user preferences with their sparse interactions. Its crucial secret lies in its ability to select samples from each user’s vast collection of unobserved items, specifically tailored to enhance the model’s optimization within its current state (Shi et al., 2023; Chen et al., 2023b). To achieve this, it necessitates the deliberate selection of negative samples characterized by (1) heightened informativeness, (2) increased discriminative capacity, and (3) enhanced precision. Such samples are typically termed as Hard Negative Samples (HNS), which are derived from their tendency to encompass more comprehensive information compared to the randomly selected samples from the corpus. Incorporating them into the training process serves to balance the positive and negative information within the dataset, thereby ensuring an unbiased optimization of the recommender. Drawing from this foundational definition of HNS, their gradient directions and magnitudes tend to vary from other random negative samples. Theoretically, more HNS can not only expedite the recommender’s convergence but also rectify the optimization direction of the global gradient, thus making it computationally possible to predict the appropriate items for users from the sparse implicit interactions (Ma et al., 2023; Zhang et al., 2021a). In the practical deployment, the inclusion of HNS enables recommenders to elevate their effectiveness and robustness simultaneously, which has been verified in many relevant studies.
@@ 2.2.1. False Negative Problem
Traditional recommenders randomly select items that users have not interacted with as negative samples (Rendle et al., 2012; Liu et al., 2021a; Zhao et al., 2018). In addition to these potentially meaningless random samples, some recommendation algorithms attempt to sample informative yet not excessively challenging negative items from the random candidates (Zhang et al., 2013; Shi et al., 2023; Chen et al., 2023b). These samples compel the recommender to delve deeper into modeling the boundary between positive and negative samples in many complex scenarios, thereby improving the recommender’s ability to discriminate among the unobserved items. The pioneering negative sampling methods have categorized these items as hard negative samples in recommendation and presented their corresponding definitions as:
Negative Samples (NS): The complete set of all items in the corpus that user has not interacted with, which serves as the fundamental candidate of all existing negative sampling strategies.
@@
The subsequent negative sampling strategies have generally focused on augmenting the ”hardness” of HNS. This includes but is not limited to introducing supplementary information (Sun et al., 2015; Wang et al., 2021b; Zhao et al., 2019; Chen et al., 2023b), enhancing the sampling weight of HNS (Zhang et al., 2013; Shi et al., 2023; Ma et al., 2023), and employing positive representation interpolation (Huang et al., 2021; Shi et al., 2022). However, the excessively high ”hardness” of NS may introduce bias during the back-propagation process, potentially leading to the poor convergence speed and quickly falling into local optimum. Furthermore, the ”hardness” of samples varies significantly across different recommendation algorithms and datasets, resulting in substantial instability in parameter selection (Zhu et al., 2022; Ding et al., 2020; Shi et al., 2023; Ma et al., 2023). To summarize, the False Negative Problem stands as a prevalent and urgent challenge to be solved in RS.
@@
| RNS (Rendle et al., 2012) | UAI | 2009 | CF | Random sampling | ||||||

### v2

@@ Abstract.
Recommender system (RS) aims to capture personalized preferences from massive user behaviors, making them pivotal in the era of information explosion. However, the presence of “information cocoons”, interaction sparsity, cold-start problem and feedback loops inherent in RS make users interact with a limited number of items. Conventional recommendation algorithms typically focus on the positive historical behaviors, while neglecting the essential role of negative feedback in user preference understanding. As a promising but easy-to-ignored area, negative sampling is proficients in revealing the genuine negative aspect inherent in user behaviors, emerging as an inescapable procedure in RS. In this survey, we first discuss existing user feedback, the critical role of negative sampling and the optimization objectives in RS and thoroughly analyze challenges that consistently impede its progress. Then, we conduct an extensive literature review on the existing negative sampling strategies in RS and classify them into five categories with their discrepant techniques. Finally, we detail the insights of the tailored negative sampling strategies in diverse RS scenarios and outline an overview of the prospective research directions toward which the community may engage and benefit.
@@ 1. Introduction
Practical RS frequently consists of more than millions of users and items, each characterized by hundreds of auxiliary attributes. Despite this, the existence of personalized preference, “Information Cocoons” and the unique feedback loops intrinsic to the RS lead to a scenario in which users only interact with a limited number of warm items, rendering the majority of items of the corpus remain consistently unfamiliar to them (Li et al., 2022; Lou et al., 2019; Chen et al., 2021c; Meng et al., 2020) (as shown in the left part of Fig.1). The core of RS lies in its ability to infer unobserved preferences from the available interaction data, with a prevailing understanding that their efficacy excels as the richness of interaction data amplifies. However, the interaction data is observational rather than experimental. Thus, the ubiquity of noise within the incomplete users’ historical interactions prevalent in real-world RS may lead to the sub-optimal decision-making in recommendation (Chen et al., 2023a; Gao et al., 2023a).
@@
In alignment with other supervised tasks, providing appropriate supervised signals (both positive and negative feedback) is indispensable in the training phase of recommendation algorithms. The absence of either of these feedback types can inevitably give rise to bias in model training (Ma et al., 2023; Zhu et al., 2022; Chen et al., 2023b). However, explicit feedback remains conspicuously absent within most recommendation datasets. Even when encompassed within datasets of such information, certain studies (Zhang et al., 2013; Ding et al., 2020; Chen et al., 2023b) have unearthed that these negative feedback indeed mirror user preferences to some extent. For instance, the behavior of assigning a low rating to a particular movie may paradoxically signify that the user has already viewed the movie, as this selection from an extensive movie pool subtly reveals the user’s underlying predilections (Ding et al., 2023; Huang et al., 2023). Consequently, the integral exploration of negative sampling strategies that are proficient in unveiling genuine negative aspects inherent in user preferences emerges as an inescapable procedure in RS.
@@
As illustrated in Fig. 1, the training stage of RS typically collect the real behaviors and select the potential negative samples (NS) via diverse sampling strategies simultaneously, and then optimize the recommender algorithm on these given positive and negative signals. The profound impact of distinct negative sampling strategies on the performance of recommendation algorithms in disparate recommendation scenarios underscores the intricacies inherent in this critical domain. Moreover, existing academic endeavors pertaining to negative sampling algorithms are diverse in RS. Varied resolutions are observed in addressing NS across distinct recommendation scenarios, indicating the absence of a universal algorithm capable of concurrently tackling all recommendation tasks (Cheng et al., 2021; Li et al., 2021; Liu et al., 2021b; Zhao et al., 2019; Shi et al., 2023; Gong et al., 2022; Zhang et al., 2013; Kim et al., 2021a; Yang et al., 2022a). Finally, the predominant strategies for tackling matching and ranking tasks revolve around the random sampling strategies (Zhao et al., 2016) and the meticulous integration of genuine negative feedback (Lian et al., 2020a; Shen et al., 2021; Gu et al., 2021) for the industrial recommender systems. Currently, there lacks a comprehensive survey that systematically categorizes and presents the existing studies. Such a survey can serve as a handbook to offer a unified exploration of the varied landscape surrounding negative sampling algorithms in RS, thereby addressing the aforementioned issues.
@@
Given the significance and popularity of RS, there are several recently published surveys also reviewed this area from different perspectives, including debiased recommendation (Chen et al., 2023a), cross-domain recommendation (Zang et al., 2022), graph-based recommendation (Gao et al., 2023c), knowledge graph-based recommendation (Guo et al., 2020b), causal inference-based recommendation (Gao et al., 2024), large language model-based recommendation (Fan, 2024; Wu et al., 2024c) and evaluation in recommendation (Bauer et al., 2023). As to negative sampling, there are surveys in other domains, including knowledge graph completion (Cai et al., 2023), deep metric learning (Kaya and Bilge, 2019), neural semantic retrieval (Guo et al., 2022; Chen et al., 2022c), image retrieval (Chen et al., 2022c), and representation learning on various types of data (Liu and Tang, 2021; Qian et al., 2020; Xu et al., 2022; Yang et al., 2020b; Chen et al., 2020a; Kamigaito and Hayashi, 2022; Wang et al., 2021a; Khoshraftar and An, 2022; Wang et al., 2017a). Compared to these research domains, negative sampling techniques employed in RS possess distinctive characteristics, necessitating a comprehensive review and summary. To bridge this gap, we review the negative sampling strategies in RS, classify them into five categories with the discrepant techniques and detail the insights of the tailored negative sampling strategies in diverse recommendation scenarios.
@@
This survey encompasses over 230 papers from the top conferences and journals within the field of recommender systems, which focus on the analysis and the design of novel negative sampling techniques. We provide a comprehensive collection of negative sampling strategies in RS from various perspectives. In particular, we search related works from the top-tier conferences on the past 20 years, including SIGIR, WWW, KDD, ACL, AAAI, IJCAI, ICDM, ICDE, RecSys, WSDM, and CIKM, as well as the leading journals such as TOIS, TKDE, TORS, TKDD, KBS, etc, with the keywords “recommendation”, “collaborative filtering”, “ranking”, “re-ranking” and “matching” in conjunction with “negative samples”, “re-weighting” and “sampling”. At the same time, to prevent omissions of relevant literature, we utilize Google Scholar, DBLP, and Arxiv to search for related studies. After getting the list of papers, we also check their references to retain the unobserved papers about negative sampling strategies. We provide Figure 2 to illustrate the statistics of our collected publications related to negative sampling in recommendation with their publication year and venue. To facilitate this survey’s role as a handbook for negative sampling techniques in RS, we will release a GitHub repository to encompass all reviewed papers along with their corresponding codes at the following link: https://github.com/hulkima/NS4RS.
@@ 2. Necessity and Challenges of Negative sampling
Negative sampling emerge as a critical and irreplaceable element in recommendation. Within this specific domain, how to effectively discover users’ real negative interests in negative sampling stands as a central research objective (Chen et al., 2023c). In this section, our first introduce two types of feedback within recommendation, then we highlight the significance of negative sampling algorithms in recommendation, with particular attention to the crucial role played by hard negative samples, and finally we analyze the existing challenges of negative sampling techniques in recommendation.
2.1. Introduction of User Feedback in Recommendation
Constrained by the exposure bias and popularity bias inherent in the real-world RS, user interaction reveals a notable unbalancedness, which constitutes a computationally insurmountable problem in recommendation (Huang et al., 2023). Regarding the available user feedback, both positive and negative feedback can reflect the user’s opinions on diverse items in the corpus, which are equally essential for modeling user preferences in RS (Zhao et al., 2018; Wu et al., 2022b). Leveraging such data to its maximum capacity can provide the recommender with a significant information gain. Following the classical studies (Wang et al., 2021c; Xu et al., 2021; Wu et al., 2022a), user feedback can be categorized into two different types as explicit feedback and implicit feedback 777Apart from explicit feedback and implicit feedback, the presence of real negative feedback is a notable factor in practical RS (Xie et al., 2021). Due to the scarcity of this type of feedback, we discuss their definitions and applications in Sec .4.4.:
Explicit Feedback: Explicit Feedback materializes in the form of precise ratings to provides unequivocal insights into user preferences (e.g., from 1.0 to 5.0).
Implicit Feedback: Implicit Feedback encompasses data derived from user behaviors without any ratings and reviews (e.g., residence time and whether to choose dislike).
From the above two types of feedback, implicit feedback is fundamentally different from the explicit one, which is captured in a structured manner (Saito et al., 2020a). Precisely, explicit feedback delivers direct insights into user preferences, while implicit feedback unravels concealed user interests. Regarding the explicit feedback, existing RS algorithms (Sun et al., 2021; Li et al., 2023; Dai et al., 2023) typically regard the user feedback with lower value (e.g., ratings of [1.0, 2.0, 3.0]) as the negative signals, which may not necessarily reflect the users’ real preferences due to the existence of selection bias in recommendation (Chen et al., 2023a). Even for the item explicitly rated as 1.0, they have been chosen by this user from an extensive candidate pool. Therefore, such negative feedback may not be as too ”negative” as they reflect this user’s preference to some extent. That is, these items appear more ”negative” than those the user is interested in, yet somewhat more “positive” when compared to items the user has never interacted with. As noted in (Wang et al., 2021c), implicit feedback can be classified into true positives, false positives, true negatives, and false negatives, among which this survey primarily concentrates on the negative aspect. For the negative implicit feedback, traditional RS methods configure all of the uninteracted items as the negative sample pool, which lacks explicitness. Meanwhile, it necessitates the recommender to elucidate users’ preferences, which may introduce long-tailed information and exposure bias in RS.
@@ 2.2. Role of Negative Sampling in Recommendation
Real-world recommender systems frequently involve more than millions of users and items, rendering the integration of all corpus into the training process prohibitively expensive (Zhang et al., 2013; Shi et al., 2023; Zhou et al., 2020). Despite the presence of a substantial number of users and items, the existence of “Information Cocoons” and the unique feedback loops within recommendation result in the majority of users interacting with only a limited number of items in the corpus (Li et al., 2022; Lou et al., 2019; Chen et al., 2021c). As a consequence, this gives rise to the well-documented issue of data sparsity in recommendations (Ma et al., 2024a, 2023; Cao et al., 2022; Xie et al., 2022b). The dynamic nature of user interests also necessitates a continuous process of adaptation and updating for recommenders (Zhao et al., 2016; Xu et al., 2021; He et al., 2016). Furthermore, the perpetual introduction of new users and items (both devoid of an adequate reservoir of historical interaction data) engenders the notorious cold start problem (Wei et al., 2021; Togashi et al., 2021).
@@
Large-scale recommendation applications typically encompass multiple phases to narrow down the candidate items for the recommender, thereby alleviating the aforementioned issues and significantly reducing their computational complexity (Huang et al., 2023; Ding et al., 2023). These algorithms generally require both positive and negative examples to model the users’ personalized preferences. Negative sampling is the critical and irreplaceable element in recommendation that could potentially improve the modeling of dynamic user preferences with their sparse interactions. Its crucial secret lies in its ability to select samples from each user’s vast collection of unobserved items, specifically tailored to enhance the model’s optimization within its current state (Shi et al., 2023; Chen et al., 2023b). To achieve this, it necessitates the deliberate selection of NS characterized by (1) heightened informativeness, (2) increased discriminative capacity, and (3) enhanced precision. Such samples are typically termed as hard negative samples (HNS), which are derived from their tendency to encompass more comprehensive information compared to the randomly selected samples from the corpus. Incorporating them into the training process serves to balance the positive and negative information within the dataset, thereby ensuring an unbiased optimization of the recommender. Drawing from this foundational definition of HNS, their gradient directions and magnitudes tend to vary from other random NS. Theoretically, more HNS can not only expedite the recommender’s convergence but also rectify the optimization direction of the global gradient, thus making it computationally possible to predict the appropriate items for users from the sparse implicit interactions (Ma et al., 2023; Zhang et al., 2021a). In the practical deployment, the inclusion of HNS enables recommenders to elevate their effectiveness and robustness simultaneously, which has been verified in many relevant studies.
2.3. Optimization Objective in Recommendation
How to design a effective and robust loss function is the universal but essential research topic in recommendation, which is bidirectionally coupled with the negative sampling techniques in recommendation to jointly establish recommenders’ optimization objective and generalization ability. Here, the loss function defines the optimization objectives for positive and negative samples, with different losses exhibiting varying degrees of dependency on and utilization of negative samples. In turn, the quality, quantity, and distribution of negative samples directly influence the gradient direction and the convergence of loss function, potentially undermining its effectiveness. Therefore, a series of studies in recommendation start to investigate effective loss functions to optimize recommenders’ convergence towards the global minima. Regarding the ratio between positive and negative samples, these approaches can be broadly categorized into One2One (one positive to one negative) and One2More (one positive to multiple negatives) schemes.
One of the most representative One2One optimization objectives in recommendation is BPR loss (Rendle et al., 2009), which assumes that user prefer the positive item over the non-observed samples. Given user , positive item and the sampled negative instance , BPR loss can be formulated as , where denotes whole training triples, denotes the Sigmoid function, and represent the score predicted by the recommender. It maximizes the score difference between positive and negative samples, which forces the recommender with relative item preferences rather than absolute scores , thereby closely aligning the goal of top-k recommendation tasks. Compared to the One2One optimization objective, One2More has attracted increasing attention due to its ability to accelerate recommender optimization and mitigate the impact of potential false negative samples on the global gradient by incorporating several negative samples (Wu et al., 2024b; Ma et al., 2023). For instance, BSL (Wu et al., 2024a) applies the same Log-Expectation-Exp structure to positive examples and multiple negatives to maintain the robustness of recommenders to the noisy positives. Adap- (Chen et al., 2023e) defines the Adaption and Fine-grained principles and utilizes Superloss to adaptively adjust the temperature by monitoring the sample losses for each user. S-DPO (Chen et al., 2024) instills ranking knowledge into the LLM via pairing multiple negatives with the positive sample to help LLM-enhanced recommenders distinguish preferred items from negatives. Benefiting from the CL paradigm, AdvInfoNCE (Zhang et al., 2023b) and PCL (Tang et al., 2024b) adaptively assigns hardness to each negative instance in an adversarial manner and collaboratively combine the absolute constraints of BCE/CE loss with CL to make the representations uniformly distributed respectively, thereby mining informative knowledge from the hard negatives. The discussion of optimization objectives in recommendation further underscores the critical role of negative sampling. We believe that the negative sampling strategies summarized in this survey can be seamlessly integrated with existing optimization objectives and will benefit from future advances in this technique to drive progress in recommendation.
@@ 2.4.1. False Negative Problem
Traditional recommenders randomly select items that users have not interacted with as negative samples (Rendle et al., 2009; Liu et al., 2021a; Zhao et al., 2018). In addition to these potentially meaningless random samples, some recommendation algorithms attempt to sample informative yet not excessively challenging negative items from the random candidates (Zhang et al., 2013; Shi et al., 2023; Chen et al., 2023b). These samples compel the recommender to delve deeper into modeling the boundary between positive and negative samples in many complex scenarios, thereby improving the recommender’s ability to discriminate among the unobserved items. Some recent negative sampling methods (Shi et al., 2023; Ma et al., 2023; Liu et al., 2023a) have explicitly distinguished these items as the HNS and further given the conception of false negative samples (FNS) in recommendation. We also leverage Fig. 3 to illustrate the distribution and the toy example of each type of samples. For instance, radios and compact discs that share the same category of “electronics” may be false negative samples (potential positive samples). Conversely, watches, tall glasses, and high-heeled shoes can serve as the negative samples with the “hardness” decreasing, which exhibit less associated with the category of “electronics”. The corresponding definitions of these samples can be presented as:
Negative Samples (NS): As discussed in Sec. 2.1, user feedback can be categorized as explicit feedback and implicit feedback. Regarding the former, NS typically denote the items which are rated with low values (e.g., [1, 2, 3]). While for the latter, NS commonly refer the complete set of all items in the corpus that user has not interacted with. This type of samples typically serves as the fundamental candidate of all existing negative sampling strategies.
@@
As highlighted in the discussion of exposure bias in (Chen et al., 2023a), user behaviors are collected from the observed data, which makes them inherently noisy, stochastic, and incomplete. However, the subsequent negative sampling strategies have generally focused on augmenting the ”hardness” of HNS. This includes but is not limited to introducing supplementary information (Sun et al., 2015; Wang et al., 2021b; Zhao et al., 2019; Chen et al., 2023b), enhancing the sampling weight of HNS (Zhang et al., 2013; Shi et al., 2023; Ma et al., 2023), and employing positive representation interpolation (Huang et al., 2021; Shi et al., 2022). In this case, certain NS considered overly ”hard” by these negative samplers might in fact represent the potential positives (Chen et al., 2023a). Therefore, incorporating them as the negative samples into the recommender optimization may introduce bias during the back-propagation process, potentially resulting in the poor convergence speed and quickly falling into local optimum and ultimately leading to the False Negative Problem in recommendation. The ”hardness” of samples varies significantly across different recommendation algorithms and datasets, resulting in substantial instability in parameter selection (Zhu et al., 2022; Ding et al., 2020; Shi et al., 2023; Ma et al., 2023) and further aggravating this problem. While some prior works (Saito, 2020; Saito et al., 2020b; Li et al., 2024; Ha et al., 2024) attempt to alleviate this phenomenon by modeling the unbiased representation with the Missing-Not-At-Random feedback, their efforts have largely centered on the theoretical guidance or the refinement of learning objectives, neglecting the more straightforward and effective strategy of negative sampling. To summarize, the false negative problem stands as a prevalent and urgent challenge of this specific research domain to be solved in RS.
@@
| RNS (Rendle et al., 2009) | UAI | 2009 | CF | Random sampling | ||||||